hey can you guys help me with this question pleasee(: i need help i gotta solve the linear systym usingg the substitution method... y=3x+2 x+2y=11
Answer by stanbon(75887)

solve the linear systym usingg the substitution method...
y=3x+2
x+2y=11
----------
Substitute for "y" and solve for "x":
x + 2(3x+2) = 11
x + 6x+4 = 11
7x = 7
x = 1
------------
Solve for "y":
y = 3x+2
y = 3*1+2
y = 5
-------------
solution (1,5)
===================
Cheers,
Stan H.
----------------